Explain the valuation formula for a constant growth stock.Explain how the formula for a zero growth stock is related to that for a constant growth stock.
1
Expert's answer
2014-12-26T09:31:40-0500
The formula for the present value of a stock with constant growth is
the estimated dividends to be paid divided by the difference between
the required rate of return and the growth rate:
P0 = (D0 ( 1+g))/(r-g)= D1/(r-g)
where
P0 = the stock price,
D0 = the current dividend,
D1 = the next dividend,
g = the dividends growth rate
r = the required return on the stock
The formula for the present value of a stock with zero growth is
dividends per period divided by the required return per period,
because If there is no growth of dividends g will be equal 0:
g=0
P0 = (D0 ( 1+0))/(r-0)= D/r